OpenStudy (anonymous):

Given the system of equations, what is the solution? 2x + 3y = 2 3x - 4y = 20

OpenStudy (lalaly):

multiply the first by 3 and the sec by 2 6x+9y=6 6x-8y=40 subtract both equations u get 17y=-34 y=-2 then substitute and find x

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

6x + 9y = 6 6x - 8y = 40 17y = -34 y = -2 x = 20 -8 / 3 = 4